OpenStudy (anonymous):

I need help with these two problems 4x^2+15x+10=0 and 2x^2+31=-15x

OpenStudy (campbell_st):

dp you need to solve them...?

OpenStudy (anonymous):

Yeah. I have NO clue on how to do it. :/

OpenStudy (campbell_st):

ok... they can't be factored so you'll need the general quadratic formula for an equation \[ax^2 + bx + c = 0\] the values of x thank make the equation true can be found using \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your 1st equation a = 4 , b = 15 and c = 10 substitute and then evaluate then 2nd equation needs to be rewritten \[2x^2 + 15x +31 = 0\] again you'll need the general quadratic formula but this time a = 2, b = 15 and c = 31 this will have complex roots hope this helps...

OpenStudy (anonymous):

Where do you go from, \[x=-15\pm \sqrt{65} \] all over 8?

OpenStudy (anonymous):

you solve for + and - both will be valid

OpenStudy (anonymous):

I don't know how to do that lol...

OpenStudy (anonymous):

make 2 equations \[x=-15+\sqrt{65}\] and \[x=-15-\sqrt{65}\]

OpenStudy (anonymous):

and just solve it?

OpenStudy (anonymous):

well if you need answer...yes =)

OpenStudy (anonymous):

what do I do about the 8. When I was writing the equation it wouldn't let me put it all over 8. but its -15 + or - sqrt of 65 all divided by 8

OpenStudy (anonymous):

ah yes divide by 8 as well as its part of the equation

OpenStudy (campbell_st):

yep....the 1st one is \[x = \frac{-15 \pm \sqrt{65}}{8}\]

OpenStudy (anonymous):

ok! and I just do the same problem (-15+ sqrt of 65 divided by 8) and thats my answer?

OpenStudy (campbell_st):

well perhaps you need it written as \[x = (-15 + \sqrt{65})/8.....and.... x = (-15-\sqrt{65})/8\]

OpenStudy (anonymous):

Yes! Idk why but that made so much more sense haha

OpenStudy (campbell_st):

I'm not sure if you need to provide and exact value... which is what has been posted or an decimal approximation... which needs to to actually calculate the value.

OpenStudy (anonymous):

Yeah it was a long decimal.

OpenStudy (campbell_st):

ok... then you need to calculate the 2 values for x...

OpenStudy (anonymous):

the second one I got -15 + and - the sq rt of -23 all over 4

OpenStudy (campbell_st):

thats right... but you will need to use i^2 = -1 to simplify the -23... this is a complex number solution to the problem and can be written \[x = \frac{-15 \pm \sqrt{23 \times -1}}{4} = \frac{-15\pm \sqrt{23 \times i^2}}{4} = \frac{-15\pm i \sqrt{23}}{4}\] hope it helps

OpenStudy (anonymous):

all of those equations kind of ran together but I think I get the idea.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (campbell_st):

well to simplify let \[i^2 = -1\] and then substitute it \[-23 = 23 \times i^2\]

OpenStudy (anonymous):

i dont think they have covered complex numbers yet - usually at that stage it is said that equation has no answer.

OpenStudy (anonymous):

yeah we haven't. idk im not good at math at all and I missed the notes over this! lol

OpenStudy (campbell_st):

well you can use the discriminant... if you know of that and say \[b^2 - 4ac = -23\] so the equation has no real roots

OpenStudy (anonymous):

ok! thanks!

OpenStudy (campbell_st):

good luck